OpenStudy (anonymous):

Given the function f(x) = log3(x + 4), find the value of f−1(3)

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

f−1(3) = 7 f−1(3) = 12 f−1(3) = 23 f−1(3) = 31

OpenStudy (anonymous):

\[\log_3(x+4)=3\] solve for \(x\)

OpenStudy (anonymous):

im lost in math :'( i dont know how to do that